OpenStudy (anonymous):

Convert 4.50 miles of Fe to atoms of Fe

Vocaloid (vocaloid):

1 mole = 6.022*10^23 atoms 4.50 moles = ?

OpenStudy (lolaismy):

Miles or moles? I think you mean moles :) 4.5 moles Fe * 6.022*10^23 atoms / 1 mole Fe = 2.77*10^24 atoms of Fe. have a good day and remember your stoichiometry ! LEARN TO LOVE THE MOLE :D